### v1

# Grasping the Essentials: Tailoring Large Language Models for Zero-Shot Relation Extraction

Abstract
Relation extraction (RE), a crucial task in NLP, aims to identify semantic relationships between entities mentioned in texts. Despite significant advancements in this field, existing models typically rely on extensive annotated data for training, which can be both costly and time-consuming to acquire. Moreover, these models often struggle to adapt to new or unseen relationships. In contrast, few-shot learning settings, which aim to reduce annotation requirements, may offer incomplete and biased supervision for understanding target relation semantics, leading to degraded and unstable performance. To provide the model with accurate and explicit descriptions of the relations types and meanwhile minimize the annotation requirements, we study the definition only zero-shot RE setting where only relation definitions expressed in natural language are used to train a RE model. Motivated by the strong synthetic data generation power of LLMs, we propose a framework REPaL which consists of three stages: (1) We utilize LLMs to generate initial seed instances based on relation definitions and an unlabeled corpora. (2) We fine-tune a bidirectional Small Language Model (SLM) using these initial seeds to learn the relations for the target domain. (3) We enhance pattern coverage and mitigate bias resulting from the limited number of initial seeds by incorporating feedback acquired from SLM’s predictions on unlabeled corpora. To accomplish this, we leverage the multi-turn conversation ability of LLMs to generate new instances in follow-up dialogues. Experiments on two datasets show REPaL achieves better zero-shot performance with large margins over baseline methods.
Grasping the Essentials: Tailoring Large Language Models for Zero-Shot Relation Extraction
Sizhe Zhou1, Yu Meng2, Bowen Jin1, Jiawei Han1 1 University of Illinois Urbana-Champaign  2 University of Virginia {sizhz, bowenj4, hanj}@illinois.edu, yumeng5@virginia.edu
Figure 1: Different types of initial seeds for low-resource RE approaches. The relation is based on P276 of FewRel Dataset (Han et al., 2018b). For instance-level seeds, we can see two instance examples fail to cover structure type entity mentions of <ENT0>.
1 Introduction
Relation extraction is an important Information Extraction (IE) task that aims to identify and classify semantic relationships between entities in natural language texts. It has wide applications in various downstream tasks including knowledge graph construction (Lin et al., 2015; Yu et al., 2020), question answering (QA) (Wang et al., 2012, 2016) and event mining (Jiao et al., 2022; Li et al., 2023c).
Existing works propose to train RE models with massive human-labeled data (Han et al., 2018a, 2020; Yamada et al., 2020). However, in real-world scenarios, it is generally challenging and costly to obtain large-scale and high-quality datasets. To tackle such data scarcity issues, recent works explore few- or zero-shot relation extraction settings. Given that pre-trained language models (PLMs) (Radford et al., 2018; Devlin et al., 2018) store factual relation knowledge (Petroni et al., 2019), the first category of methods proposes to directly prompt them with human-crafted patterns (Zhou et al., 2023a; Sainz et al., 2021; Chia et al., 2022; Han et al., 2022; Chen et al., 2022a). Another category of methods fine-tunes PLMs to capture the relation semantics from different types of texts (Gao et al., 2020; Zhao et al., 2023).
More recently, the remarkable reasoning and generation power of LLMs make it possible to attain better RE performance with fewer human annotations: Some methods exploit the in-context learning paradigm or assistance from other language models to enhance semantic understanding of target relations (Wan et al., 2023; Ma et al., 2023). Other works conduct relational data synthesis to mitigate noise or hallucination issues (Chia et al., 2022; Li et al., 2023b; Sun et al., 2024).
Despite the unprecedented advance achieved by PLM/LLM-based few-shot or zero-shot relation extraction works, they suffer from two issues. The first issue is the underutilization of relation definitions, which offer explicit and accurate relation semantics and are less biased than a limited number of demonstrations. Specifically, many few-shot or zero-shot RE approaches start with instance-level seeds such as relation instances and relation triples (see Fig. 1). Though simple and versatile, such seeds occasionally result in partial coverage of target relation semantics, thus the resulting RE models may fail to grasp the complete relational semantics. Unlike other tasks such as text classification where label semantics are relatively simple (Meng et al., 2020), relation semantics are generally directional and multifaceted which involve entity-entity interactions and entity-related requirements (see Fig. 1). Such relation semantic complexity requires detailed elaborations and is considered in the annotation process of most RE datasets  [^1] For example, SemEval-2010 Task 8 dataset guideline and TAC KBP guidelines but frequently gets neglected in RE model developments.
Another issue is that the potential of LLMs for zero-shot relation extraction is still far from being fully exploited. First, while LLMs are primarily developed and extensively pre-trained for domain adaptation (Radford et al., 2019), current LLM-based zero-shot RE approaches often struggle to achieve satisfactory performance. This suggests that the zero-shot learning capabilities of LLMs may not have been fully utilized. Second, most LLMs are designed to perform multi-turn conversations and excel in seeking feedback from the dialogue history. Such a feature has shown great potential in a vast range of other NLP tasks including knowledge-intensive or complex reasoning question-answering tasks (Trivedi et al., 2022; Zhou et al., 2023b, c). Nevertheless, zero-shot and few-shot LLM-based RE approaches mainly perform the task with single-turn usages of LLMs.
In this work, we focus on definition-based low-resource relation extraction. We study a new zero-shot RE setting where only relation definitions, instead of relation instances, are provided. In addition to the fact that relation definitions serve as less biased initial seeds, such a task setup is realistic as: (1) many downstream applications such as QA tasks already have explicit definitions of interested relations and obtaining such supervision is generally straightforward; and (2) such a setting highlights the importance for RE systems to continuously adapt to new relation types based on corresponding definitions without maintaining a large amount of seen instances and re-training models.
To better leverage LLMs for the proposed zero-shot RE setting, we propose a novel zero-shot RE framework, REPaL. Firstly, REPaL prompts LLMs to generate positive instances conditioned on the definition and samples negative instances from the unlabeled corpus, thus constructing an initial training set. We then train a bidirectional SLM on these instances for training and inference efficiency.
Secondly, to improve pattern coverage of instance generation and rectify the bias within the generated instances and the trained SLM, REPaL acquires instance-level feedback from the trained SLM’s inference results on the unlabeled corpus. Such feedback is utilized to: (1) leverage LLMs’ multi-turn conversational ability to recognize the pattern coverage and existing LLM-SLM’s bias, and then generate instances with new positive patterns and rectified positive patterns, and (2) leverage LLMs’ reasoning ability to diagnose the SLM’s bias and further generate targeted negative instances to rectify such bias. The whole framework performs iterative refinements in which more and better relation instances are accumulated to improve the task-specialized RE model. To foster reproducibility, our data and codes will be organized and available soon.
Our contributions are as follows:
  • We propose the definition only zero-shot RE setting along with modified evaluation datasets to address the deficiency of traditional low-resource RE approaches.
  • We propose a novel zero-shot RE framework that leverages LLMs and unlabeled corpus to generate relation instances and self-improves the semantic coverage while rectifying bias by automatically acquiring and reasoning over feedback.
  • Extensive quantitative and qualitative experiments demonstrate the effectiveness as well as the potential of our proposed task setup and framework.
2 Problem Formulation
Definition 2.1. Definition Only Zero-Shot Relation Extraction. Specifically, we assume that, for any target relation $r(E_{0},E_{1})\in\mathcal{R}(E_{0},E_{1})$  [^2] This work focuses on sentence-level binary relation extraction setting, so two entity mentions are considered for each instance., only one associated relation definition $d(E_{0},E_{1})$ is given. Here $\mathcal{R}(E_{0},E_{1})$ denotes the whole binary relation space and $d(E_{0},E_{1})$ can be a single sentence or a document specifying the target relation $r(E_{0},E_{1})$. $E_{0}$ and $E_{1}$ are two entity placeholders.
The goal of Definition Only Zero-Shot Relation Extraction task is to extract all relation instances that belong to target relation $r(E_{0},E_{1})$ from a given set of relation instances $\{(s^{i},e_{0}^{i},e_{1}^{i})\}_{i=1}^{n}$. Here $s^{i}$ denotes the $i$-th context text, while $e_{0}^{i}$ and $e_{1}^{i}$ denote two entity mentions inside $s^{i}$. Compared with the traditional zero-shot RE task setting, our task assumes no seen data but only target relation definitions. Our task also assumes the unknown negative relation space while mainstream zero-shot RE assumes known information of all test relations. This makes our task more challenging, but aligns better with real-world scenarios.
3 Methodology
Figure 2: REPaL framework. The trained SLM-Based RE Model is used in inference stage.
Our proposed REPaL is an iterative framework which consists of three major components: (1) Definition-based seed construction with LLM and the unlabeled corpus (Sec. 3.1); (2) Pattern learning with RE-specialized SLM (Sec. 3.2); and (3) Feedback-driven instance improvement and bias rectification (Sec. 3.3). The overview of REPaL can be found in Fig. 2.
3.1 Definition-Based Seed Construction with LLM and Unlabeled Corpus
LLMs, such as the GPT family, are pre-trained for the domain adaptation ability (Radford et al., 2019). They have demonstrated to contain factual relation knowledge (Petroni et al., 2019) and are widely used as data generator for various tasks such as text classification (Meng et al., 2022) and dialogue systems (Abdullin et al., 2024). Follow-up evaluation studies have shown that LLMs are relatively skilled at constrained content generation, storytelling, and rationale generation (Sun et al., 2023; Keskar et al., 2019). Following such observations as well as the goal to tackle the data scarcity issue of zero-shot RE, REPaL first leverages LLMs to generate $K_{p_{0}}$ positive seeds, $\{(s_{p}^{i},e_{0,p}^{i},e_{1,p}^{i})\}_{i=1}^{K_{p_{0}}}$, conditioned on the target relation definition $r(E_{0},E_{1})$. We carefully design three prompts for this step (shown in Appendix A) to encourage the LLM to generate comprehensive patterns from three perspectives in terms of complexity: brief, medium, and implicit. Based on exploratory experiments, they yield different types of relation patterns in accordance with our design expectations. Details of experiments and quantitatively how choice of different prompts affect the results are in Sec. 5.3.
In addition to positive seed generation, constructing negative seeds is also vital for better extraction for the target relation $r(E_{0},E_{1})$, as our task setting assumes no prior knowledge on the negative relation space in order to mimic the real-life deployments (Li et al., 2022). Our negative seed construction is based on the hypothesis that, with a normal large-scale unlabeled corpus, the proportion of target relation instances is relatively small. Thus REPaL conducts random sampling over the unlabeled corpus, which is expected to yield an acceptable set of negative instances denoted as $\{(s_{n}^{i},e_{0,n}^{i},e_{1,n}^{i})\}_{i=1}^{K_{n_{0}}}$. In case of extreme cases, REPaL further designs counter-measures in Sec. 3.3.
3.2 Pattern Learning with RE-Specialized SLM
Given the fact that LLMs are computationally expensive to be fine-tuned, while pure in-context learning (ICL) is hindered by LLM’s sequence length limit, LLM’s capacity to comprehend demonstrations, and LLM’s temporal and monetary expense (Ma et al., 2023), we leverage SLMs as tunable and task-specialized extractors. Specifically, we adapt a natural language inference (NLI) model (Obamuyide and Vlachos, 2018; Liu et al., 2019; Sainz et al., 2021) to be relation-specific binary classifiers for simplicity and leave other architectures for future research.
For each training instance $(s^{j},e_{0}^{j},e_{1}^{j})$, the input is formulated as in NLI style:
$\texttt{Premise}_{j}:$ $=s^{j},$
$\texttt{Hypothesis}_{j}:$ $=d(E_{0}=e_{0}^{j},E_{1}=e_{1}^{j}).$
Given a SLM model $\mathcal{M}$, we obtain the encoded sequence hidden states $\mathbf{H}$ by:
$\mathbf{H}=\mathcal{M}(\texttt{Premise}_{j}\,\texttt{[SEP][SEP]}\,\texttt{Hypothesis}_{j})$
and the NLI logits $\mathbf{z}=[z_{E},z_{N},z_{C}]\in\mathbb{R}^{3}$ is computed as:
$\mathbf{z}=\mathbf{W}\cdot\mathbf{H}_{[CLS]}+\mathbf{b}.$
Finally, $P_{j}$, the probability of instance $(s^{j},e_{0}^{j},e_{1}^{j})$ following relation $r(E_{0},E_{1})$, is computed as the normalized logit of ENTAILMENT label:
$P_{j}=\frac{e^{{z}_{E}}}{\sum_{c\in\{C,N,E\}}e^{{z}_{c}}},$
where $C,N,E$ denote logits $\mathbf{z}$’s indices for NLI label CONTRADICTION, NEUTRAL, ENTAILMENT respectively. And the binary classification loss (Shannon, 1948) for relation $r(E_{0},E_{1})$ is defined as:
$\begin{split}\mathcal{L}=-\frac{1}{|B|}\sum_{\begin{subarray}{c}(s^{j},e_{0}^{j},e_{1}^{j})\in B\end{subarray}}&[y_{j}\log(P_{j})\\ +&(1-y_{j})\log(1-P_{j})],\end{split}$ (1)
where $B$ denotes the batched data sampled from $\{(s_{p}^{i},e_{0,p}^{i},e_{1,p}^{i})\}_{i=1}^{K_{p_{0}}}\bigcup\{(s_{n}^{i},e_{0,n}^{i},e_{1,n}^{i})\}_{i=1}^{K_{n_{0}}}$.
3.3 Feedback-Driven Instance Improvement and Bias Rectification
After obtaining the trained SLM relation extractor, REPaL further introduces a follow-up feedback-driven iterative refinement approach with intuitions listed as follows: (1) the initially generated positive seeds might only have a partial relation pattern coverage. Instead of one-time generation, iterative follow-up generations by incorporating feedback from examining previously generated results are better for pattern coverage. (2) bias could exist after the previous two steps (Sec. 3.1 and 3.2). One potential source of bias is the generated positive seeds or the randomly sampled negative seeds. Another potential source of bias is the training process of SLMs over limited training samples. Identifying and rectifying bias requires a feedback-based follow-up refinement approach.
Specifically, we first obtain the inference results on the unlabeled corpus leveraging trained SLM RE model. The feedback is constructed by sampling instances within certain prediction confidence range. Feedback is constructed for follow-up positive instance generation and for follow-up negative instance generation respectively. The only difference between the above two feedback constructions lies in two distinct prediction confidence ranges.
Secondly, for follow-up positive instance generation, we fully leverage the multi-turn conversational feature of LLMs where the positive instance generation history is maintained and a new turn of conversation asking for positive instances with different relation patterns is appended. The feedback is integrated into the new turn of conversation so that LLMs can identify the correct and incorrect patterns learned by SLMs.
For follow-up negative instance generation, we divide it into two sub-steps: the first sub-step is feedback-driven negative relation definition generation and the subsequent sub-step is negative instance generation based on the negative relation definitions. LLM is queried to examine the constructed feedback and identify the incorrect patterns learned by SLM so that newly generated negative relations can address such patterns. We also query LLM to generate near-miss negative relation definitions to better distinguish hard cases. After acquiring the negative relation definitions, the second sub-step is the same as the initial positive seed generation process in Sec. 3.1. The prompts used in this section are placed in Appendix B. After obtaining all the follow-up positive and negative instances, we iteratively train the SLM RE model.
4 Experiments
4.1 Evaluation Dataset Construction
Dataset #test rels #test #unlabel
DefOn-FewRel 14 9,798 100,000
DefOn-ReTACRED 11 2,148 100,000
Table 1: Brief statistics about datasets constructed.
Model DefOn-FewRel DefOn-ReTACRED
[…]
Fully-Supervised
RoBERTa NLI 77.34 99.06 85.89 77.61 98.35 86.03
Few-Shot
GPT-3.5 ICL 46.65 72.86 54.97 40.28 64.55 48.72
Zero-Shot
Random Guess 7.14 49.76 12.49 9.10 50.02 15.37
GPT-3.5 50.79 69.76 54.23 49.58 38.48 42.52
RoBERTa NLI 39.67 94.34 49.40 26.94 97.55 40.64
ZS-BERT* 41.00 40.51 40.73 20.20 17.70 18.81
RelationPrompt* 75.18 66.08 70.34 51.67 51.26 51.40
RelationPrompt (NoGen)* 0.00 0.00 0.00 2.98 0.83 1.25
RE-Matching* 76.47 71.80 74.05 54.43 50.22 52.16
REPaL (Ours) 78.19 82.25 77.93 68.47 80.52 68.42
Table 2: Evaluation results of relation extraction models under zero-shot and few-shot settings. Zero-shot models marked by * are trained on seen relation instances and require all negative test relations to be known. REPaL is trained on 15p15n initial instances and further trained with additional 15p15n follow-up instances.
We first construct datasets for model evaluation on the Definition Only Zero-Shot Relation Extraction task. Our constructed datasets are based on FewRel (Han et al., 2018b) and ReTACRED (Zhang et al., 2017; Stoica et al., 2021) respectively. As FewRel’s relations are from WikiData’s  [^3] WikiData main page properties, we query the property definitions and slightly modify them to become complete sentences with entity placeholders [ENT0] and [ENT1] (corresponding to $E_{0}$ and $E_{1}$ as in definition notation $d(E_{0},E_{1})$). As TACRED/ReTACRED’s data are based on 2009-2014 TAC KBP (Zhang et al., 2017), we annotate relation definitions based on the guidelines (especially the slot definitions  [^4] TAC KBP guidelines) as well as related adjustments introduced in ReTACRED paper (Stoica et al., 2021).
The evaluation relations are chosen based on the following general principles: (1) The selected relation should has abundant instances for evaluation (generally, we set a minimum requirement of 200 instances); (2) The annotated instances should follow the officially documented relation definitions and the sentence should be sufficient to deduce the target relations without external knowledge; and (3) The selected relations should be semantically disjoint without ambiguity. These principles lead to the brief dataset statistics in Table 1.
To construct unlabeled corpus, we down-sample the FewRel unlabeled corpora (from WikiData) (Han et al., 2018b; Gao et al., 2020) to 10,000 instances. For ReTACRED (Zhang et al., 2017; Stoica et al., 2021), our constructed unlabeled corpora includes all unselected instances and sampled instances from FewRel unlabeled corpus so that the unlabeled corpus size also reaches 10,000. More details on the two constructed datasets can be found in Appendix C.
4.2 Baseline Methods
We compare our method with the following baselines under Fully-Supervised, Few-Shot, and Zero-Shot settings: (1) Random Guess: gives uniform random binary predictions. (2) GPT-3.5 (Brown et al., 2020; Ouyang et al., 2022): uses gpt-3.5-turbo-0125 model and formulates RE as a double choice problem, answering whether two entity mentions in each test instance follow the given target relation definition. GPT-3.5 ICL on DefOn-ReTACRED uses randomly sampled 10 positive and 10 negative (abbreviated as 10p10n) gold instances as demonstrations while adopts 5p5n for DefOn-FewRel (better than 10p10n and 2p2n). (3) RoBERTa NLI (Devlin et al., 2018; Liu et al., 2019): our adopted SLM RE backbone model (Sec. 3.2). It uses 100p100n for DefOn-FewRel and 70p70n for DefOn-ReTACRED for Fully-Supervised setting. (4) ZS-BERT (Chen and Li, 2021): a Siamese-based zero-shot RE framework trained on seen labeled instances and gives prediction by nearest neighbour search comparing encoded sentence representations and relation definition representations. (5) RelationPrompt (Chia et al., 2022): a Seq2Seq-based zero-shot RE framework that trains GPT-2 (Radford et al., 2019) to generate relation instances conditioned on relation names and train BART (Lewis et al., 2020) to generate the extracted relation triple on seen data. For unseen data, it fine-tune BART on synthetic data generated by trained GPT and then predicts. (6) RE-Matching (Zhao et al., 2023): a Siamese-based zero-shot RE model that encodes entity types and relation definitions for each relation on seen data and conducts nearest neighbour search for prediction on unseen data. We also test on RE-Matching (NoGen) that removes the data generator.
Note that ZS-BERT, RelationPrompt and RE-Matching require being trained on labeled seen data and their predictions on unseen data are conducted in a multi-way classification manner where information of all negative relations is leveraged. These three baselines are trained on 61 FewRel relations with another 5 FewRel relations in dev set (each relation has 700 instances).
4.3 Experiment Setup
Evaluation Methods.
Following our Definition Only Zero-Shot Relation Extraction setting, evaluation is conducted similar to cross validation where each iteration one test relation is treated as positive target relation and remaining test relations serve as negative test relations (our setting assumes all negative relations are unknown). Macro $F_{1}$ is computed by averaging across all iterations. Reported results in this work are the average over random seeds 1, 2, and 3. See Appendix E for implementation details.
5 Results and Analysis
5.1 Main Results
The main results are shown in Table 2. REPaL achieves much better performance compared with all few-shot and zero-shot baselines with large margins. As the strong baselines are pre-trained on FewRel data, not ReTACRED data, REPaL’s advance is much larger in DefOn-ReTACRED. This shows the effectiveness of our method in low-resource settings and the robustness across different benchmarks as we do not rely on any benchmark-specific supervision. Compared with fully-supervised baselines, it shows there is still room for improvement. This is likely due to our task’s assumption on unknown negative test relation space which is normally avoided in mainstream zero-shot RE works.
5.2 Ablation Study
Model DefOn-FewRel
[…]
REPaL
   RoBERTa NLI 39.67 94.34 49.40
   w/ P_init 67.22 90.11 73.76
  +P_feedback 72.56 87.79 76.84
   +N_feedback 78.19 82.25 77.93
Table 3: Ablation results with 15p15n as initial setting. P_init, P_feedback, and N_feedback denote initial positive generation, feedback-driven follow-up positive generation and negative generation respectively.
To investigate the effectiveness of our framework design, we conduct ablation studies with results shown in Table 3. The initial seed generation brings considerable performance advance which shows LLM’s power of domain adaptation is fully leveraged given a clear relation definition. Based on experiments on Sec. 5.4, the best $F_{1}$ achieved by a single seed among all initial seed generation setups is 75.92 (p=60, n=60, random seed=3). So by adopting our feedback-driven follow-up generation design, in total p=30 & n=30 surpasses this upper bound with approximately 2 points.
Furthermore, the negative follow-up instance generation further boosts the precision score, demonstrating its effectiveness on rectifying SLM’s bias in distinguishing positive and negative relations.
5.3 Effect of Positive Templates
Model DefOn-FewRel
[…]
REPaL
 w/ medium 67.93 86.40 72.77
    +brief 68.57 88.52 74.38
     +implicit 68.25 86.85 73.42
Table 4: Evaluation results w/o follow-up generation and conditioned only on different initial positive generation templates. 30p30n instances are gathered for all settings.
Table 5 in Appendix A shows the adopted initial positive seed construction templates. Note that no ICL is adopted in REPaL’s generation step. Analysis in Appendix A yields the conclusion that the generated instance’s patterns generally follow the prompt instructions well, covering instances with brief, medium-length, and implicit patterns respectively.
The quantitative results with different combinations of positive instance generation prompts are shown in Table 4. We can see that results are generally robust against different prompt combinations. And leveraging medium and brief prompts yields the best result. It’s likely because two evaluation datasets have short or medium length/complexity instances. However, in our main experiments, all three prompts are adopted to cover diverse types of generated patterns.
5.4 Effect of More Initial Seeds
Figure 3: $F_{1}$ (%) scores for different setups in REPaL’s first step of Definition-Based Seed Construction with LLM and Unlabeled Corpus.
As recent LLMs have enabled larger context windows, a naïve method for improving framework performance is to directly query LLMs for more generated instances. So we conduct such trials to investigate whether more initially generated seeds can bring more benefits. The quantitative results are shown in Fig. 5, Fig. 6, and Fig. 3. We can see that synthetic data by LLM is generally beneficial, but generating more initial seeds does not guarantee better results. Larger p&n or larger n:p ratio both lead to higher precision and lower recall.
One explanation for these trends is that more positive seeds does not bring larger positive pattern coverage but results in over-fitting. However, more randomly sampled negative seeds lead to larger negative pattern coverage (Examining distant labels of original FewRel unlabeled corpora shows #potential relations are 744). So the model tends to give positive predictions on instances with some dominant positive patterns but identify other minority patterns as negative. Also note that when n&p are small, usually recall is pretty high while precision is low. This may also be caused by our task assumption on unknown negative relation space. The above observations and analysis further justify our design of feedback-driven follow-up generation and refinement paradigm.
5.5 Derive Relation Definition from Few-Shot Instances
Figure 4: Micro $F_{1}$ (%) score of model trained with few-shot instances (Trained w Few-Shot Ex) and model trained with instances from our relation definition derivation and instance generation approach (Def Deduce+Ex Gen). The error bar/band denotes averaged value $\pm$ standard deviation.
In this section, we show that the commonly adopted RE few-shot setup can be converted to our definition-oriented setup while achieving much better results. We take KnowPrompt (Chen et al., 2022b) as the underlying N-way classification model where N equals the number of test relations. It’s a prompt tuning RE model with robust and strong few-shot performance and does not require hand-crafted prompts.
We experiment on two groups: (1) Trained w Few-Shot Ex (baseline group) which has KnowPrompt trained on few-shot instances for evaluation, and (2) Def Deduce+Ex Gen which uses GPT-4 to derive each relation’s definition given 4-shot instances, generate 15 new instances based on the definition, and train KnowPrompt for evaluation.
The LLM derived relation definitions compared with original relation definitions are shown in Appendix F.1 and the experiment results are shown in Fig. 4 and Fig. 7 (in Appendix F.2). The derived relation definitions show that LLM is capable of deducing suitable yet generalizable relation definitions based on few-shot instances. However, the coverage of derived definitions is limited by the coverage of few-shot instances. This is in accordance with our motivation for definition-based low-resource RE setup.
Fig. 4 and Fig. 7 show that our definition derivation and instance generation approach achieves much better performance than the model trained only on few-shot instances. This indicates the approach extends the relation patterns conveyed by the few-shot instances. However, we can see our (15 generated + 4 gold shots) trained model has slightly lower performance than 16 gold shots trained model which is likely due to the partial coverage of relation semantics conveyed by the 4 gold shots instances. This further illustrates the importance of capturing actual relation definitions instead of few-shot data for low-resource RE approaches.
6 Related Works
Zero-Shot Relation Extraction
Our work is related to zero-shot RE (Levy et al., 2017). Majority zero-shot RE approaches mainly leverage clustering, label-verbalization, or Siamese-based architectures (Rahimi and Surdeanu, 2023; Chen and Li, 2021; Chia et al., 2022; Li et al., 2023a) which seek for the instance-instance similarity or the similarity between the relation instances and the unseen relations’ information. Chen and Li (2021) utilize relation descriptions for zero-shot RE but their approach still relies on seen data to align relation descriptions with instances in a supervised manner. Li et al. (2023b) adopt the relation descriptions but only for verifying synthesized data with the instance-level seeds. LLM-based RE works focus on designing prompting strategies or LLM alignment to tackle zero-shot RE (Li et al., 2023a; Zhang et al., 2023; Wei et al., 2023). Our work is distinguished from pre-LLM zero-shot RE works as they heavily rely on the supervision from massive seen data and the complete negative relation space. And the majority do not focus on relation definitions. Our work is different from LLM-based zero-shot RE works as we emphasize both the rich relation definitions for data synthesis and synergy between SLM and LLM.
Definition-Driven Text Mining
Another related line of research is definition-driven text mining. BERTNet (Hao et al., 2023) applies definitions for distilling entities from LM parametric knowledge. Label definitions/descriptions have also been proven to be powerful in text classification (Gao et al., 2023). In zero-shot RE, several PLM-based works have utilized relation definitions (Chen and Li, 2021; Zhao et al., 2023) but they mainly focus on computing instance-definition similarities. In our work, LLM is used to distill patterns described in definitions and extend or rectify the learning of SLM starting from definitions.
7 Conclusion
In this work, we have introduced a new zero-shot RE task where only relation definitions instead of seen-unseen relation instances are provided. Correspondingly, we have proposed REPaL which leverages LLMs and unlabeled corpus to generate relation instances and iteratively self-improves the generation pattern coverage while rectifying the bias by automatically acquiring and reasoning over sampled feedback. Quantitative experiments and qualitative analysis on our two modified datasets show the effectiveness and robustness of our framework as well as our large-margin advance over SOTA baselines. Further exploratory experiments show that generating more data in a single-turn conversation does not yield larger pattern coverage. We also proposed a derive-definition-and-then-generate approach which achieves much better performance than just utilizing few-shot instances. This gives insights into low-resource RE works to capture the complete relation semantics to avoid partial coverage by few-shot instances.
Limitations
In this work, we mainly experimented on GPT-3.5 and GPT-4 as their instruction following performance is competent so that we do not need to introduce in-context learning in most of the time. So one follow-up work is to explore other LLMs to see their generation capability compared to the GPT series. Another limitation of this work is on the evaluation RE datasets. The scale of the evaluation datasets could be extended given enough budgets for computations or LLM API calling. And other datasets could be explored although RE still lacks large scale yet high quality datasets. Thirdly, prompt engineering and hyperparameter search are not conducted. For the sake of better performance in downstream tasks, future works could compensate this.
Ethics Statements
Since our goal is to solve sentence-level RE tasks where the text contexts are sufficient to derive the relation, factualness of the relation triples is not a strict requirement or a vital factor for the training instances. Therefore, in generative data synthesis, we do not further verify the factualness of the generation results and we simply count on the GPT-4 model. So follow-up works could explore this and other related approaches should also be careful if they want to adapt our work to downstream tasks necessitating factualness such as factual question answering.
Acknowledgements
Research was supported in part by US DARPA KAIROS Program No. FA8750-19-2-1004 and INCAS Program No. HR001121C0165, National Science Foundation IIS-19-56151, and the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No. 2118329. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of DARPA or the U.S. Government.
Appendix A Prompt Templates Used in Definition-Based Seed Construction with LLM
  Prompt Name Prompt Template Example Generated Instances (Relation=P106)
  brief
A binary relation between entity placeholders <ENT0> and <ENT1> is defined by “\${Relation Definition}”. Under sentence-level relation extraction setting, generate \${Number of Examples} examples (numbered from 1 to \${Number of Examples}) expressing the same relation, where <ENT0> is replaced with actual entity mention and is prefixed with tag <ENT0> and suffixed with tag </ENT0> , <ENT1> is replaced with actual entity mention and is prefixed with tag <ENT1> and suffixed with </ENT1> . Do not overfit the pattern of the definition. Try as many different relation patterns or relation expressions as possible.
1. Before becoming a renowned author, <ENT0>Agatha Christie</ENT0> had a career as a <ENT1>nurse</ENT1> during the World War I period.
2. The award-winning film director <ENT0>Quentin Tarantino</ENT0> once worked at a video rental store, which heavily influenced his future as a <ENT1>filmmaker</ENT1>.
3. Known for his groundbreaking theories in physics, <ENT0>Albert Einstein</ENT0> held the position of a <ENT1>patent clerk</ENT1> in his early career, which gave him the time to develop his early theories.
medium
A binary relation between entity placeholders <ENT0> and <ENT1> is defined by “\${Relation Definition}”. Under sentence-level relation extraction setting, generate \${Number of Examples} examples (numbered from 1 to \${Number of Examples}) expressing the same relation, where <ENT0> is replaced with actual entity mention and is prefixed with tag <ENT0> and suffixed with tag </ENT0> , <ENT1> is replaced with actual entity mention and is prefixed with tag <ENT1> and suffixed with </ENT1> . Other content requirements: 1. Do not overfit the pattern of definition. Try as many different relation patterns or relation expressions as possible. 2. Generate rich and informative related contexts before and after each entity.
1. During the 1990s, the renowned physicist <ENT0>Stephen Hawking</ENT0> dedicated his life to unraveling the mysteries of the universe, a commitment that made him synonymous with the role of <ENT1>theoretical physicist</ENT1>.
2. Before stepping into the political arena, <ENT0>Barack Obama</ENT0> honed his skills and passion for change as a <ENT1>community organizer</ENT1> in Chicago, working tirelessly on initiatives that aimed to improve living conditions for the city’s residents.
3. In her illustrious career spanning several decades, <ENT0>Marie Curie</ENT0> shattered gender barriers and set new benchmarks in scientific research, earning worldwide acclaim as a pioneering <ENT1>chemist</ENT1> and physicist.
implicit
A binary relation between entity placeholders <ENT0> and <ENT1> is defined by “\${Relation Definition}”. Under sentence-level relation extraction setting, generate \${Number of Examples} examples (numbered from 1 to \${Number of Examples}) expressing the same relation, where <ENT0> is replaced with actual entity mention and is prefixed with tag <ENT0> and suffixed with tag </ENT0> , <ENT1> is replaced with actual entity mention and is prefixed with tag <ENT1> and suffixed with </ENT1> . Other content requirements: 1. Do not overfit the pattern of definition. Try as many different relation patterns or relation expressions as possible. 2. Generate rich and informative related contexts before and after each entity. 3. The relation patterns or relation expressions should be implicit or complicated
1. Born into a family with a history of academic excellence, it was no surprise when <ENT0>Albert Einstein</ENT0> ventured into the realms of theoretical physics, a field where he made his mark by unraveling the mysteries of the universe, thus becoming synonymous with the profession that <ENT1>theoretical physicist</ENT1> would forever be associated with his name.
2. The battlefields of ancient Macedonia were graced by the strategic genius of <ENT0>Alexander the Great</ENT0>, whose military campaigns across the vast expanses of the known world were unmatched, firmly establishing him not just as a king but as one of the greatest <ENT1>military leaders</ENT1> history has ever witnessed, shaping the course of the civilizations that came after him.
3. In the annals of crime-solving history, <ENT0>Sherlock Holmes</ENT0> stands out as a fictional detective who, through his acute observation and deductive reasoning, set the gold standard for the <ENT1>detective</ENT1> profession, influencing countless others in the field of criminal investigation despite his creator, Sir Arthur Conan Doyle’s, initial reluctance to make him a permanent fixture in literary history.
Table 5: Prompt templates used in Definition-Based Seed Construction with LLM (Sec. 3.1). Words in blue denote the placeholders for filling in contents indicated by their surface names.
Table 5 contains the three prompt templates used for generating initial positive seeds using LLMs. Example instances generated by corresponding prompts are also shown inside the table. Our goal of designing such prompts is to cover all the patterns for target relations. From the generated example instances, we can see that the pattern complexity (or more naïvely, the sentence length) exhibits significant distinctions, especially between implicit and other two prompts. The pattern complexity (or more naïvely, the sentence length) well follows the instructions conveyed by each type of prompt and well represents the prompt name, brief, medium, and implicit correspondingly.
Appendix B Prompt Templates Used in Feedback-Driven Instance improvement and Bias Rectification
  Prompt Name Prompt Template
  Follow-up Positive Instance Generation
[…]
Follow-up Negative Relation Definition Generation
[…]
Table 6: Prompt templates used in Feedback-Driven Instance Improvement and Bias Rectification (Sec. 3.3). Words in blue denote the placeholder for filling in contents indicated by their surface names.
Table 6 contains prompt templates for our feedback-driven follow-up positive instance generation and negative relation definition generation respectively. After obtaining the negative relation definitions, we simply leverage the medium instance generation template in Table 5 to generate negative relation instances. We take this template as our purpose of negative instance generation is to rectify the existing bias instead of pursuing complete negative relation pattern coverage. Furthermore, Table 4 demonstrates that the performance difference between the usages of different templates is minor.
Appendix C Details of Constructed DefOn-FewRel and DefOn-ReTACRED Datasets
  Original Dataset Relation Label : Definition Frequency
  FewRel
P106: <ENT1> was/is the occupation of <ENT0> (a person)
700
FewRel
P1344: <ENT0> (a person or organization) was/is a participant of <ENT1> (an event)
700
FewRel
P136: <ENT1> was/is the genre or the field of work of <ENT0> (a creative work or an artist)
700
FewRel
P1411: <ENT1> was/is the award nomination received by <ENT0> (a person, organisation, or creative work)
698
FewRel
P241: <ENT1> was/is the military branch to which <ENT0> (a military unit, award, office, or person) belonged/belongs
700
FewRel
P26: <ENT1> was/is the married spouse (husband, wife, partner, etc.) of <ENT0>
700
FewRel
P276: <ENT1> was/is the location of <ENT0> (an object, structure or event)
700
FewRel
P3373: <ENT1> and <ENT0> had/have at least one common parent (<ENT1> is the sibling, brother, sister, etc. including half-sibling of <ENT0>)
700
FewRel
P40: <ENT1> was/is the child (not stepchild) of <ENT0>
700
FewRel
P400: <ENT1> was/is the platform or platform version for which <ENT0> (a work or a software product) was/is developed or released
700
FewRel
P410: <ENT1> was/is the military rank achieved by or associated with <ENT0> (a person or a position)
700
FewRel
P57: <ENT1> was/is the director(s) of <ENT0> (a film, TV-series, stageplay, video game or similar)
700
FewRel
P84: <ENT1> was/is the architect or architectural firm responsible for designing <ENT0> (a building)
700
FewRel
P974: <ENT1> was/is the watercourse that flowed/flows into <ENT0> (a watercourse)
700
Table 7: DefOn-FewRel relation labels, definitions, and corresponding instance frequencies.
  Original Dataset Relation Label : Definition Frequency
  ReTACRED
org:founded: <ENT1> was/is the date/time on which <ENT0> (an organization) was founded
148
ReTACRED
org:founded_by: <ENT0> (an organization) was/is founded by <ENT1> (a person, organization, or geopolitical entity)
200
ReTACRED
org:member_of: <ENT0> (an organization) was/is a member (or part) of <ENT1> (an organization)
200
ReTACRED
org:members: <ENT1> (an organization) was/is a member (or part) of <ENT0> (an organization)
200
ReTACRED
per:age: <ENT1> was/is the (reported, approximate, or previous) age of <ENT0> (a person)
200
ReTACRED
per:cause_of_death: <ENT1> was/is the explicit cause of death for <ENT0> (a person)
200
ReTACRED
per:charges: <ENT0> (a person) was/is charged with, alleged to have been involved in, or convicted of <ENT1>
200
ReTACRED
per:children: <ENT1> (a person) was/is the child (biological child, adopted child, or step-child) of <ENT0> (a person)
200
ReTACRED
per:date_of_death: <ENT1> was/is the date/time of <ENT0> (a person)’s death
200
ReTACRED
per:parents: <ENT1> (a person) was/is the parent (biological parent, step-parent, or adoptive parent) of <ENT0> (a person)
200
ReTACRED
per:spouse: <ENT1> (a person) was/is the spouse of <ENT0> (a person)
200
Table 8: DefOn-ReTACRED relation labels, definitions, and corresponding instance frequencies.
Table 7 and Table 8 show the relation labels, constructed definitions and corresponding frequencies of our modified DefOn-FewRel and DefOn-ReTACRED.
To get quality evaluation datasets, we first conduct data cleaning with the requirements as: (1) The two entity mentions should not overlap; (2) The entity mentions should not be pronouns such as I, he, and she. For the construction of test sets for DefOn-FewRel and DefOn-ReTACRED, we follow the principles listed in Sec. 4.1 to select relations from FewRel (Han et al., 2018b) train and dev sets and ReTACRED dataset (Zhang et al., 2017; Stoica et al., 2021). Since ReTACRED dataset is not balanced, we generally require that test relations have at least 200 instances for the evaluation effectiveness.
For the construction of unlabeled corpus, we globally set up the size of unlabeled corpus as 10,000 as a trade-off between the corpus scale and the computation costs. Furthermore, we require that selected unlabeled samples should not be repeated. Namely, for any two unlabeled samples, the sentence, the head entity mention and the tail entity mention can not all be the same. As FewRel already comes with an unlabeled corpora, we first gather all the unlabeled samples with similar entity mention pairs  [^5] Here we use fuzzy string match and sets the similarity threshold as 90 out of 100. which results in 9,806 unlabeled instances. The remaining 90,194 unlabeled instances are obtained by random sampling. As ReTACRED does not have an unlabeled corpora, we down-sample each test relation’s instances to 200 instances. The exceeding test instances and all the other unselected relations’ instances are placed inside the unlabeled corpus, which results in 64,992 instances. We then obtain 91 instances with similar entity mention pairs. The remaining 34,917 unlabeled instances are composed of 24,442 instances that are most similar to test set instances  [^6] We adopt all-mpnet-base-v2 checkpoint of Sentence-BERT (Reimers and Gurevych, 2019) and gauge the similarity in terms of cosine similarity of sentence embeddings. and 10,475 instances randomly sampled (the two numbers come from the 7:3 ratio setup).
Appendix D Effect of More Initial Seeds
Figure 5: Precision (%) scores for different set-ups in REPaL’s first step of Definition-Based Seed Construction with LLM and Unlabeled Corpus.
Figure 6: Recall (%) scores for different set-ups in REPaL’s first step of Definition-Based Seed Construction with LLM and Unlabeled Corpus.
The precision (%) scores and recall (%) scores discussed in Sec. 5.4 are shown in Fig. 5 and Fig. 6 respectively.
Appendix E Implementation Details
E.1 Baselines
In consideration of OpenAI API calling expense, the GPT-3.5 baseline results are from the evaluation over down-sampled test sets (30 down-sampled test instances for each relation). For all the other baselines, we adopted their original hyperparameters reported in corresponding publications.
E.2 REPaL
The LLM model we use is gpt-4-0125-preview with max tokens set to 4,096, presence penalty set to 0.15 and temperature set to 1.0. The OpenAI API cost to reproduce one random seed’s results of REPaL in Table 2 is around 2 dollars (far cheaper than direct inference with GPT-3.5 on test sets). The SLM checkpoint used is roberta-large-mnli. All model training is conducted on 2 or 4 NVIDIA RTX A6000 GPUs.
For the step of Definition-Based Seed Construction with LLM and Unlabeled Corpus, we follow the templates in Table 5. For the results reported in Table 2, we choose the setting of 15 initial positive seeds and 15 initial negative seeds based on the analysis in Sec. 5.4. As we have three types of prompts for positive seeds generation, each prompt will contribute to generating 5 initial positive seeds. For the step of Pattern Learning with RE-Specialized SLM, we train the SLM with 5 epochs using AdamW optimizer (Loshchilov and Hutter, 2017) with learning rate equal to 3e-5 and batch size equal to 64.
For the step of Feedback-Drive Instance Improvement and Bias Rectification, we first leverage the trained SLM to conduct inference on the unlabeled corpus after which each unlabeled instance will be associated with a score as the probability of being positive. Then, for follow-up positive instance generation, we conduct random sampling from all instance with score higher than 0.85 as we want the sampled feedback instances to reflect the model’s learning outcome for the target positive relation. For follow-up negative relation definition generation, we conduct random sampling from all instances with score higher than 0.50 as we want to see both the confident predictions and less confident predictions to identify the existing bias. Note that continuing from the initial positive seed generation, there will be three threads of dialogue history for follow-up positive instance generation corresponding to three types of prompts in Table 5. So for each thread of dialogue, we fill in different groups of sampled feedback instances to maximize the feedback coverage. For both the follow-up positive instance generation and the follow-up negative instance generation, the number of feedback instances for each prompt input is set to 10. For follow-up negative relation definition generation, we set the number of generated negative relation definitions to be 5 and the number of total follow-up negative instances to be 15. After obtaining the feedback-driven follow-up instances, we continue to train the SLM RE model for 3 epochs with all the other hyperparameters the same as our previous SLM training step.
Appendix F LLM-Based Relation Definition Derivation
  Prompt Name Prompt Template
  Few-Shot Definition Derivation
[…]
Train Instance Generation
A binary relation between entity placeholders <ENT0> and <ENT1> is defined by “\${Derived Relation Definition}”. Under sentence-level relation extraction setting, generate additional \${Number Of Additional Examples to Generate} examples (numbered from 1 to \${Number Of Additional Examples to Generate} expressing the same relation, where <ENT0> is replaced with actual entity mention and is prefixed with tag <ENT0> and suffixed with tag </ENT0> , <ENT1> is replaced with actual entity mention and is prefixed with tag <ENT1> and suffixed with </ENT1> . \${Gold Few-Shot Examples for ICL} Do not overfit the pattern of the definition. Try as many different relation patterns or relation expressions as possible.
Table 9: Prompt template used in deriving original relation definitions given few-shot relation instances and generating new relation instances based on the derived relation definition and gold few-shot instances (Sec. 5.5). Words in blue denote the placeholder for filling in contents indicated by their surface names.
Our adopted prompt template for deriving relation definitions based on few-shot instances is shown in Table 9. Note that we leveraged a fixed 3 relation definition demonstrations for in-context learning across all relations so that the LLM can give the relation definition in our desired format for automatic parsing. After getting the relation definition, we use the prompt template in Table 9 to generate 15 instances for each derived relation. Note that the instance generation prompt is basically the same as brief prompt in Table 5 except that it integrates the gold few-shot instances as in-context learning demonstrations.
F.1 Relation Definitions Derived by LLM From Few-Shot Instances
  Gold Definition Gold Few-Shot Instances For Derivation Derived Definition
<ENT1> was/is the occupation of <ENT0> (a person)
[…]
2. WWF Hall of Famer Bob Backlund and Extreme Championship Wrestling <ENT1>manager</ENT1> <ENT0>Bill Alfonso</ENT0> also made surprise appearances during the event .
3. In May 2010 , Paratici moved from Sampdoria to Juventus , along with Director General Giuseppe Marotta and <ENT1>Manager</ENT1> <ENT0>Luigi Delneri</ENT0> .
[…]
<ENT1> is the profession in which <ENT0> (a person) works or has worked.
<ENT0> (a person or organization) was/is a participant of <ENT1> (an event)
1. He only saw limited action in <ENT1>Euro 2000</ENT1> as cover for left - back <ENT0>Arthur Numan</ENT0> .
2. <ENT0>Francesco Cameli</ENT0> was a sailor from Italy , who represented his country at the <ENT1>1928 Summer Olympics</ENT1> in Amsterdam , Netherlands .
3. <ENT0>Giannin Andreossi</ENT0> ( born July 2 , 1902 , date of death unknown ) was a Swiss ice hockey player who competed in the <ENT1>1928 Winter Olympics</ENT1> .
[…]
<ENT1> is the major international sports competition in which <ENT0> (an athlete) has competed.
<ENT1> was/is the genre or the field of work of <ENT0> (a creative work or an artist)
1. Another version , dating from c. 1616 , was given in c. 1790 to <ENT0>Joshua Reynolds</ENT0> by the Duke of Leeds in exchange for a Reynolds self - <ENT1>portrait</ENT1> .
2. Teixeira is a former member of indie rock bands Ik Mux and Boris Ex - Machina , as well as the <ENT1>hip hop</ENT1> group <ENT0>Da Weasel</ENT0> and industrial metal band Bizarra Locomotiva .
3. Beautiful Stories for Ugly ChildrenMUSHROOMHEAD To Release ’ Beautiful Stories For Ugly Children ’ In September is the seventh studio album by <ENT1>industrial metal</ENT1> band <ENT0>Mushroomhead</ENT0> .
4. Wales is portrayed in the 1976 <ENT1>western film</ENT1> “ <ENT0>The Outlaw Josey Wales</ENT0> ” by actor and director Clint Eastwood .
<ENT1> is the genre or type of art (music, painting, film) associated with <ENT0> (an artist, band, or cultural artifact).
<ENT1> was/is the award nomination received by <ENT0> (a person, organisation, or creative work)
1. On January 24 , 2012 , he was nominated for an <ENT1>Academy Award for Best Adapted Screenplay</ENT1> for the movie “ <ENT0>Moneyball</ENT0> ” .
2. “ <ENT0>The Great Santini</ENT0> ” received two Academy Award nominations : <ENT1>Best Actor in a Leading Role</ENT1> ( Duvall ) and Best Actor in a Supporting Role ( O’Keefe ) .
3. “ <ENT0>Born This Way</ENT0> ” ( 2011 ) , Gaga ’s second studio album , accrued three nominations at the 54th Annual Grammy Awards , including her third consecutive nomination for <ENT1>Album of the Year</ENT1> .
4. As a producer , he has been nominated for <ENT1>Best Picture</ENT1> for three other films : “ Raging Bull ” , “ <ENT0>The Right Stuff</ENT0> ” , and “ Goodfellas ” .
<ENT1> is the award category for which <ENT0> (films, albums, or individuals associated with entertainment productions) has been nominated.
<ENT1> was/is the military branch to which <ENT0> (a military unit, award, office, or person) belonged/belongs
[…]
2. Lieutenant - Colonel <ENT0>Gordon Graham Donaldson</ENT0> was a senior officer in the <ENT1>British Army</ENT1> who died as a result of illness contracted during the disastrous Walcheren Campaign in 1809 .
3. <ENT0>Raphael Semmes</ENT0> was an officer in the <ENT1>United States Navy</ENT1> from 1826 to 1860 and the Confederate States Navy from 1860 to 1865 .
[…]
<ENT1> is the military organization (such as an army or navy) with which <ENT0> (an individual, specified by their role or rank) has served or been associated.
<ENT1> was/is the married spouse (husband, wife, partner, etc.) of <ENT0>
1. The film is about <ENT0>Carolyn Cassady</ENT0> ’s recollection of life with husband <ENT1>Neal Cassady</ENT1> and Jack Kerouac , and her concern that the truth about these men is being lost in their mythos .
2. Maximilian married Duchess Helene in Bavaria , daughter of <ENT1>Duke Maximilian Joseph in Bavaria</ENT1> and his wife <ENT0>Princess Ludovika of Bavaria</ENT0> , on 24 August 1858 at Possenhofen Castle .
3. In 1916 his younger daughter , <ENT0>Nadejda</ENT0> ( “ Nada ” ) married <ENT1>Prince George of Battenberg</ENT1> , older son of Prince Louis by Queen Victoria ’s granddaughter , Princess Victoria of Hesse - Darmstadt .
4. The fourth and youngest son of King <ENT1>John II of France</ENT1> and his wife , <ENT0>Bonne of Luxembourg</ENT0> , Philip was the founder of the Burgundian branch of the House of Valois .
<ENT1> is the spouse or partner of <ENT0> (an individual), indicating a marital, romantic, or partnership connection between the two entities.
<ENT1> was/is the location of <ENT0> (an object, structure or event)
1. At the <ENT0>2014 Winter Olympics</ENT0> , Hudec won the bronze medal in the super - G at <ENT1>Rosa Khutor</ENT1> .
2. On the night of 22 January 1942 during the <ENT0>Battle of the Points</ENT0> , Japanese troops of the 16th Division attempted a landing on the west coast of southern <ENT1>Bataan</ENT1> .
3. Since the Netherlands did boycott the Moscow Olympic Games Brasser represented his National Olympic Committee at the <ENT0>1980 Summer Olympics</ENT0> in <ENT1>Tallinn</ENT1> , USSR under the Dutch NOC flag .
4. The bridge Norrbro stretches past the Riksdag on <ENT0>Helgeandsholmen</ENT0> and further south to <ENT1>Stockholm Old Town</ENT1> and the Royal Palace .
<ENT1> is the location or venue where <ENT0> (an event such as sports competitions, battles, or significant historical or cultural events) took place or was hosted.
<ENT1> and <ENT0> had/have at least one common parent (<ENT1> is the sibling, brother, sister, etc. including half-sibling of <ENT0>)
1. Together they had three sons : Antonio , <ENT1>Arturo</ENT1> , and <ENT0>Alejandro</ENT0> .
[…]
3. <ENT1>Arjuna</ENT1> was the fourth one to fall after Draupadi , <ENT0>Sahadeva</ENT0> and Nakula .
[…]
<ENT1> is the sibling, specifically the brother, of <ENT0>.
Table 10: Comparison between gold relation definitions and few-shot (4-shot) derived relation definitions (random seed=1).
  Gold Definition Gold Few-Shot Instances For Derivation Derived Definition
<ENT1> was/is the child (not stepchild) of <ENT0>
1. He was the son of Flemish painter <ENT1>Jan Massys , Matsys , or Metsys</ENT1> and the grandson and namesake of <ENT0>Quentin Massys or Metsys</ENT0> .
2. She married <ENT1>Lu Jing</ENT1> , who was born to <ENT0>Lu Kang</ENT0> and another daughter of Zhang Cheng ; both Sun He ’s daughter and Lu Jing therefore were Zhang Cheng ’s maternal grandchildren .
3. She is the wife of Bollywood actor , <ENT0>Jackie Shroff</ENT0> and mother of <ENT1>Tiger Shroff</ENT1> and Krishna Shroff .
4. His uncle was polymath <ENT0>Lionel Penrose</ENT0> , whose children include mathematician <ENT1>Oliver Penrose</ENT1> , polymath Sir Roger Penrose , chess grandmaster Jonathan Penrose , and geneticist Shirley Hodgson .
<ENT1> is a direct family member (such as a son, grandson, wife, or mother) of <ENT0>, specified by their familial relationship.
<ENT1> was/is the platform or platform version for which <ENT0> (a work or a software product) was/is developed or released
1. The <ENT1>NES</ENT1> version of <ENT0>Shadowgate</ENT0> also carries the distinction of being one of the few NES games to be available in a Swedish language version .
2. In case of incidents <ENT0>Plumbr</ENT0> provides its users with information on problem severity , problem ’s root cause location in source code or <ENT1>JVM</ENT1> configuration and lists steps needed to take to remediate the problem .
3. In 2013 , “ <ENT0>Mega Man Xtreme</ENT0> ” was made available on the Virtual Console of Japan ’s Nintendo eShop for the <ENT1>Nintendo 3DS</ENT1> .
4. Prior to <ENT0>Windows 2000</ENT0> , Windows NT ( and thus PE ) supported the MIPS , Alpha , and <ENT1>PowerPC</ENT1> ISAs .
<ENT1> is the platform, console, or environment for which <ENT0> (software applications, games, or operating systems) is designed or available.
<ENT1> was/is the military rank achieved by or associated with <ENT0> (a person or a position)
1. The son of Robert Langton Douglas , he was a half - brother to <ENT1>Marshal of the Royal Air Force</ENT1> <ENT0>William Sholto Douglas , 1st Baron Douglas of Kirtleside</ENT0> .
2. <ENT0>Dwight Edward Aultman</ENT0> , <ENT1>Brigadier General</ENT1> , United States Army .
3. He then served in the 27th U - boat Flotilla along with “ <ENT1>Korvettenkapit0̆0e4n</ENT1> ” <ENT0>Erich Topp</ENT0> .
4. <ENT0>Axel Schimpf</ENT0> ( born 1 October 1952 ) is a retired “ <ENT1>Vizeadmiral</ENT1> ” ( vice admiral ) of the German Navy .
<ENT1> is the military rank of <ENT0> (a military personnel).
<ENT1> was/is the director(s) of <ENT0> (a film, TV-series, stageplay, video game or similar)
1. Cummins ’s photographs have been used extensively in cinema and TV documentaries including <ENT1>Grant Gee</ENT1> ’s <ENT0>Joy Division</ENT0> and John Dower ’s .
2. In 2014 , Zhang starred in <ENT1>Tsui Hark</ENT1> ’s wuxia film “ <ENT0>The Taking of Tiger Mountain</ENT0> ” .
3. Starting her career in 2005 , she acted in the Malayalam film “ <ENT0>Boyy Friennd</ENT0> ” directed by <ENT1>Vinayan</ENT1> .
4. Kaif had her first success in Bollywood when she appeared opposite Salman Khan in <ENT1>David Dhawan</ENT1> ’s romantic comedy “ <ENT0>Maine Pyaar Kyun Kiya ?</ENT0> ” .
<ENT1> is the director who directed the movie <ENT0>.
<ENT1> was/is the architect or architectural firm responsible for designing <ENT0> (a building)
[…]
2. Tampere Cathedral ( Lars Sonck , 1900 ) , <ENT0>National Museum</ENT0> , Helsinki ( <ENT1>Herman Gesellius , Armas Lindgren and Eliel Saarinen</ENT1> , 1902 ) .
3. Its designer was <ENT1>George Gilbert Scott</ENT1> , <ENT0>Busbridge Church</ENT0> \2013 Church of England Official gateway to the church .
4. He served a seven - year apprenticeship with <ENT1>Sir Charles Barry</ENT1> , the architect of the <ENT0>Houses of Parliament</ENT0> and Manchester Art Gallery .
<ENT1> is the architect or group of architects who designed <ENT0> (a building or architectural project).
<ENT1> was/is the watercourse that flowed/flows into <ENT0> (a watercourse)
1. The <ENT1>Cerchez River</ENT1> is a tributary of the <ENT0>Ceair River</ENT0> in Romania .
2. The lake flows into the <ENT1>River Mangfall</ENT1> , a tributary of the <ENT0>River Inn</ENT0> and thence the River Danube .
3. The <ENT1>Veljul Mic River</ENT1> is a tributary of the <ENT0>Veljul Mare River</ENT0> in Romania
4. A small part of the district along the eastern boundary drains into the east - flowing <ENT1>River Loud</ENT1> , a tributary of the <ENT0>Hodder</ENT0> .
<ENT1> is a tributary of the <ENT0> (rivers or water bodies).
Table 11: (Continued from Table 10) Comparison between gold relation definitions and few-shot (4-shot) derived relation definitions (random seed=1).
Table 10 and Table 11 show the LLM derived relation definitions based on the gold 4-shot instances. The table also contains the ground truth relation definitions for reference. We can see that for most of the FewRel relations, LLM successfully recovers the gold relation definitions. The derived definitions also reveal that one major difficulty is to specify the entity type constraints as few-shot instances may only convey a partial set of entity types which misguides LLMs to deduce a partial entity type constraints in the derived relation definitions.
F.2 Macro F1 Scores of Few-Shot Method against Definition-Based Method
Figure 7: Macro $F_{1}$ (%) scores of model trained with few-shot instances (Trained w Few-Shot Ex) and model trained with instances from our relation definition derivation and instance generation approach (Def Deduce+Ex Gen). The error bar/band denotes averaged value $\pm$ standard deviation.
The macro $F_{1}$ scores of the experiments conducted in Sec. 5.5 are shown in Table 7. Since the DefOn-FewRel dataset is almost balanced, the micro $F_{1}$ and macro $F_{1}$ are close. So we put macro $F_{1}$ here for reference.
Appendix G Case Study
Figure 8: Example interaction dialogue which demonstrates the initial seed generation and feedback-driven follow-up positive instance generation.
Figure 9: Example interaction dialogue which demonstrates the feedback-driven generation of negative relation definitions.
In order to showcase how LLMs function in accordance with our design expectations, we provide some dialogue cases in Fig. 8 and Fig. 9. From the example dialogues, we can see that GPT-4 is considerably capable of identifying and summarizing the existing patterns in the dialogue history. The conversational feature also improves the interpretability of the results which can be further exploited with human-in-the-loop methods for future research or downstream applications.
We can also see that feedback-driven negative relation definition generation is of great potential. As the feedback instances may contain false predictions which can be directly taken by LLM to generate new negative relations and effectively recitify the SLM’s bias. Such feature directly boost the precision score as shown in Sec. 5.2. Hence, we can see that our design expectations are well fulfilled and this also qualitatively explains the performance boost of our proposed model.
Appendix H Error Analysis
  Example Target Relation Majority False Positive Predicted Relations Example Instance of False Positive Predicted Relations
  P40: <ENT1> was/is the child (not stepchild) of <ENT0>
1. P26 (218): <ENT1> was/is the married spouse (husband, wife, partner, etc.) of <ENT0>
2. P3373 (165): <ENT1> and <ENT0> had/have at least one common parent (<ENT1> is the sibling, brother, sister, etc. including half-sibling of <ENT0>)
1. <ENT1> Lord Louis Mountbatten </ENT1> ( Hugh Bonneville ) arrives at Viceroy ’s House in Delhi in 1947 with his strong - willed wife <ENT0> Edwina </ENT0> ( Gillian Anderson ) and daughter Pamela . (Gold: P26 | Pred: P40 | Pos Prob: 0.995)
2. Daughter of <ENT1> Sancho IV </ENT1> and of <ENT0> María de Molina </ENT0> , Infanta Beatrice was born in Toro . (Gold: P26 | Pred: P40 | Pos Prob: 0.998)
3. <ENT1> Ruby Aldridge </ENT1> is the daughter of former Playboy playmate Laura Lyons and artist and graphic designer Alan Aldridge , and younger sister of fashion model <ENT0> Lily Aldridge </ENT0> . (Gold: P3373 | Pred: P40 | Pos Prob: 0.959)
4. <ENT0> Jan van Hoogstraten </ENT0> ( 1628–1654 ) was a Dutch Golden Age painter who died young in Vienna , where he was living with his older brother <ENT1> Samuel van Hoogstraten </ENT1> . (Gold: P3373 | Pred: P40 | Pos Prob: 0.998)
P410: <ENT1> was/is the military rank achieved by or associated with <ENT0> (a person or a position)
P241 (376): <ENT1> was/is the military branch to which <ENT0> (a military unit, award, office, or person) belonged/belongs
1. In November 1966 , retired <ENT1> USMC </ENT1> Major <ENT0> Donald Keyhoe </ENT0> and Richard H. Hall , both of NICAP , briefed the panel . (Gold: P241 | Pred: P410 | Pos Prob: 0.996)
2. General Sir <ENT0> Hew Whitefoord Dalrymple </ENT0> , 1st Baronet ( 3 December 1750 – 9 April 1830 ) was a <ENT1> British Army </ENT1> general and Governor of Gibraltar . (Gold: P241 | Pred: P410 | Pos Prob: 0.995)
P57: <ENT1> was/is the director(s) of <ENT0> (a film, TV-series, stageplay, video game or similar)
P84 (131): <ENT1> was/is the architect or architectural firm responsible for designing <ENT0> (a building)
1. The <ENT0> Yokosuka Arts Theatre </ENT0> , part of the Bay Square complex by <ENT1> Kenzo Tange </ENT1> , is a venue for opera , orchestral concerts , chamber music , and films . (Gold: P84 | Pred: P57 | Pos Prob: 0.981)
2. Memorials such as <ENT1> Maya Lin </ENT1> ’s <ENT0> Vietnam Veterans Memorial </ENT0> and Peter Eisenman ’s Memorial to the Murdered Jews of Europe are also said to reflect themes of trace and erasure . (Gold: P84 | Pred: P57 | Pos Prob: 0.979)
Table 12: Error analysis of the predictions made by SLM-based RE model. The contents in red denote the number of false positive predictions for a specific relation. The contents in blue denote the prediction details made by SLM-based RE model. Gold refers to the gold relation label of an instance. Pred refers to the predictions made by our model. Pos Porb means the predicted probability of the instance following the target relation.
Table 12 shows the major source of false positive predictions of the final tuned SLM RE model, it can be seen that the majority of false positive predictions are concentrated on a few similar negative relations. As our proposed task setting assumes the unknown negative relation space, one challenge appears to be learning the positive relation against the unknown and infinitely many negative relations. To address such challenge, our model derive targeted negative relations based on the feedback of model inference. Based on the results in Table 12, we can see that the challenge is not fully eliminated which serves as a promising future research direction. Another feature seen from the false positive predicted instances is that some typical false positive instances actually express the target relation in addition to its gold relation. However, the target relation is not expressed by the tagged entity mention pair. This may indicate that better RE architectures which well model the position awareness of target entity pairs can be adopted for improving the overall performance. We also leave this for future studies.